# Detecting Toxic Flow

### v1

@@ 6 Trade Prediction Effectiveness and Missed Opportunities
[…]
Figure 14: Histogram of predicted probability of a toxic trade. Toxicity horizon is ten seconds. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022.
The predictions of PULSE are distributed sparsely. This is a desirable feature when the broker wishes to find an optimal cutoff probability to internalise-externalise trades. For example, the concentration of the predictions we see for MLE in the top right panel of Figure 14 explains the abrupt changes we see in Figure 12 and Figure 13 for MLE, LogR, and RF.
7 Robustness analysis
7.1 Inventory aversion
In this section, we study how the value of inventory aversion parameter $\Phi$ affects the results of the AUC and the internalisation-externalisation strategy. Intuitively, as the value of inventory aversion parameter increases, the broker is willing to increase the cutoff probability to internalise a trade if it would reduce the absolute value of her inventory; similarly, she is willing to decrease the cutoff probability to internalise a trade if internalising the trade increases the absolute value of her inventory.
A higher value of the inventory aversion parameter does not necessarily mean that the broker externalises more trades, it means that the broker is keener to externalise trades that decrease the absolute value of her inventory and less keen to internalise trades that increase the absolute value of her inventory. Figure 15 (left) shows the internalised volume as a function of the cutoff probability $\mathfrak{p}$ and the inventory aversion parameter $\Phi$. Figure 15 (right) shows the PnL and avoided loss of the internalisation-externalisation strategy in (5.2) as a function of the inventory aversion parameter $\Phi$ for a toxicity horizon of 60s using PULSE.
Figure 15: Left panel: percentage of internalised volume as a function of the inventory aversion parameter $\Phi$. Right panel: PnL and avoided loss of the internalisation-externalisation strategy as a function of the inventory aversion parameter $\Phi$.
As the value of the inventory aversion parameter $\Phi$ increases, roughly, the broker internalises less of the order flow. On the other hand, the performance of the internalisation-externalisation strategy is stable for small values of the value of the inventory aversion parameter. When the value of the inventory aversion parameter becomes large, the strategy performs poorly.
Figure 16 shows the daily volume of the internalisation-externalisation strategy in (5.2) as a function of the inventory aversion parameter $\Phi$ for a toxicity horizon of 60s using PULSE. This is a detailed version of what the left panel of Figure 15 describes. Indeed, the higher the value of the inventory aversion parameter, the less daily internalised volume.
7.2 Performance for independent clocks
In Section 2.3 we construct 168 of the 175 features using three clocks: transaction-clock, time-clock, and volume-clock. Here, we explore the AUC of our models when we use only one of the three clocks to construct the features, we omit MLE because its predictions are feature-free. Recall that 168 out of the 183 features used in the calculations are measurements of 8 variables with three clocks and seven time horizons (168=3×8×7168=3$\times$ 8$\times$ 7168 = 3 × 8 × 7, see Subsection 2.3). Here, instead, we evaluate the model using 56 clock-features instead of 168, that is, we use either (i) the transaction-clock (txn), (ii) the time-clock (time), or (iii) the volume-clock (vol). Table 3 shows the AUCs, where we observe that the models are robust to the choice of clock.
@@ 7.2 Performance for independent clocks
Table 3: AUC for toxicity horizon of 30s by model and by clock. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022.
We conclude that for PULSE, considering all three clocks does not add value. For RF on the other hand, the extra feature-engineering exercise adds value.
7.3 The added value of having one model per client
@@
Figure 17: AUC outperformance over no-trader id

### v2

@@ C.1 Inventory aversion
C.2 Performance for independent clocks
In Subsection 3.2 we constructed 168 of the 175 features with three clocks: transaction-clock, time-clock, and volume-clock. Here, we explore the AUC of our models when we use only one of the three clocks to construct the features, we omit MLE because its predictions are feature-free. Recall that 168 out of the 183 features used in the calculations are measurements of 8 variables with three clocks and seven time horizons ($168=3\times 8\times 7$, see Subsection 3.2). Here, instead, we evaluate the model with 56 clock-features instead of 168, that is, we use either (i) the transaction-clock, (ii) the time-clock, or (iii) the volume-clock. Table 3 shows the AUCs, where we observe that the models are robust to the choice of clock.
@@ C.2 Performance for independent clocks
Table 3: AUC for toxicity horizon of 30s by model and by clock. EUR/USD currency pair over the period 1 August 2022 to 21 October 2022. Here, time is the time-clock, txn is the transaction-clock, and vol is the volume clock.
We conclude that for PULSE, considering all three clocks does not add value. For RF on the other hand, the extra feature-engineering exercise does add value.
C.3 The added value of one model per client
@@
Figure 18: Ability to predict toxic flow. Difference between AUC measure without and with trader identification.